If you read 15 minutes per day every day and end up reading 12 books of 200 pages each in one year, what is your reading rate in pages per minute?
Answer by nerdybill(7384)

If you read 15 minutes per day every day and end up reading 12 books of 200 pages each in one year, what is your reading rate in pages per minute?
.
Pages read:
12*200 = 2400 pages
.
Minutes reading:
365*15 = 5475 minutes
.
That's 2400 pages per 5475 minutes
or
2400/5475 = 0.44 page per minute